The formula for final velocity of an object starting from zero and undergoing constant acceleration is v = 1/2 * a * t ^ 2 What is the time it will take a car accelerating 4 m/s ^ 2 ? 50 m/s


1
Expert's answer
2021-04-26T15:56:31-0400

The formula for final velocity of an object starting from zero and undergoing constant 

acceleration is "V=\frac{1}{2}×a×t"2

where a is acceleration, t is time, v is velocity.

It follows from this eq uation that:

"t^2=\frac{2V}{a}"


"t=\sqrt\frac{2V}{a}"

Substitute v = 50 (m/s), a = 4 (m/s2) into the previous formula:

"t=\sqrt\frac{2×50}{4}"

"=\sqrt{25}=5seconds"


Answer: It will take 5 seconds